### v1

@@ 4.2 Explicitly Surface-aligned and Adaptive Gaussians
Before determining the Gaussian covariances, we introduce a local coordinate frame as illustrated in Fig. 3(a). This frame with origin at $\boldsymbol{v}_{1}$ comprises axes $\boldsymbol{t}_{i}=\tilde{\boldsymbol{t}}_{i}/\|\tilde{\boldsymbol{t}}_{i}\|,\ i\in\{1,2,3\}$, with:
$\tilde{\boldsymbol{t}}_{1}=\boldsymbol{a}\times\boldsymbol{b},\quad\tilde{\boldsymbol{t}}_{2}=\boldsymbol{a},\quad\tilde{\boldsymbol{t}}_{3}=(\boldsymbol{a}\times\boldsymbol{b})\times\boldsymbol{a}$ (7)
@@
If the angles of a triangle are all close to $\pi/3$, the covariance matrix in the local coordinate frame can be naturally defined as $\boldsymbol{\Sigma}_{e}=\text{diag}(\epsilon,r^{2},r^{2})$, representing a thin layer of Gaussian surrounding a small area of the surface. The choice of $r$ is detailed in the appendix. For triangles with a more irregular aspect ratio, where the area is challenging to approximate by a disc, it is beneficial to linearly transform the Gaussian distribution with covariance $\boldsymbol{\Sigma}_{e}$. Fig. 3(b) visualizes this transformation within the local coordinate frame. Here, the current triangle is denoted as $\boldsymbol{T}=[\boldsymbol{v}^{\prime}_{1},\boldsymbol{v}^{\prime}_{2},\boldsymbol{v}^{\prime}_{3}]$, and the reference equilateral triangle as $\boldsymbol{E}=[\boldsymbol{v}^{\prime}_{1},\boldsymbol{v}^{\prime}_{2},\hat{\boldsymbol{v}}^{\prime}_{3}]$, where $\boldsymbol{v}^{\prime}_{i}$ represents the local coordinates corresponding to $\boldsymbol{v}_{i}$, and $\hat{\boldsymbol{v}}^{\prime}_{3}=l*[0,\frac{1}{2},\frac{\sqrt{3}}{2}]^{T}$, with $l=\|\boldsymbol{v}_{2}-\boldsymbol{v}_{1}\|$.
By solving the equation $\boldsymbol{M}\boldsymbol{E}=\boldsymbol{T}$, we obtain:
$\boldsymbol{M}=\begin{bmatrix}1&0&0\\ 0&1&\frac{2\boldsymbol{v}^{\prime(2)}_{3}-l}{\sqrt{3}l}\\ 0&0&\frac{2\boldsymbol{v}^{\prime(3)}_{3}}{\sqrt{3}l}\\ \end{bmatrix},$ (8)
where $\boldsymbol{v}^{\prime(i)}_{3}$ denotes the $i$-th component of $\boldsymbol{v}^{\prime}_{3}$. The covariance matrix for the face Gaussians in world space is then given by:
@@
4.3 Gaussian Appearance Model
There are no direct connections between Gaussians derived from $(\mathcal{V}_{t},\mathcal{F}_{t})$ and those from $(\mathcal{V}_{t-1},\mathcal{F}_{t-1})$, where $t$ denotes the $t$-th training step. Hence, we are unable to optimize a constant group of per-Gaussian colors.
A straightforward approach would be to compute colors from neighboring Gaussians at the last step, formulated as $\boldsymbol{c}_{i,t}=\frac{1}{|\mathcal{N}|}\sum_{j\in\mathcal{N}}\boldsymbol{c}_{j,t-1}$, where $\boldsymbol{c}_{i,t}$ denotes the color of the $i$-th Gaussian at step $t$, and $\mathcal{N}$ is the list of neighbors based on the Gaussian center. However, this method may require significant computation due to the neighbor search between two large sets of Gaussians. Additionally, averaging colors within a local region might lead to an over-smoothing issue.
To address these challenges, we introduce a volumetric appearance model, $\mathcal{A}:\mathbb{R}^{3}\mapsto\mathbb{R}^{D}$, with $D$ representing the color space dimension. At each step, we determine the color as $\boldsymbol{c}=\mathcal{A}(\boldsymbol{\mu})$ for the Gaussian centered at $\boldsymbol{\mu}$. To enhance efficiency and achieve a compact and continuous representation, we implement $\mathcal{A}$ using hash-grid positional encoding [18] and a lightweight MLP, inspired by NvdiffRec [19]. This appearance model can be robustly learned, even as the mesh changes from $(\mathcal{V}_{t-1},\mathcal{F}_{t-1})$ to $(\mathcal{V}_{t},\mathcal{F}_{t})$.
@@ 4.3 Gaussian Appearance Model
4.4 Rendering and Optimization
As the Gaussians we construct retain the same form as those in the original 3DGS, we employ the same method defined in Eq. (2) for image rendering. Our method relies only on photometric supervision and utilizes the loss function:
@@ 4.4 Rendering and Optimization
Upon the convergence of the mesh and appearance model, we fix the topology defined by faces $\mathcal{F}$ and proceed to jointly refine the vertices and the associated 3D Gaussians. During this phase, Gaussian centers continue to be derived from the face vertices using established barycentric coordinates. However, for covariance and color, we adopt a different approach since we can optimize a consistent set of Gaussian attributes during this phase. We sample the spherical harmonics (SHs) from the appearance model $\mathcal{A}$ and make them a learnable, per-Gaussian variable as in the original 3DGS. Moreover, the Gaussian covariance is parameterized into learnable 2D scaling $\hat{\boldsymbol{s}}\in\mathbb{R}^{2}$ and a complex number $\boldsymbol{a}i+\boldsymbol{b}$ representing rotation, which are similar to the process in SuGaR [10]. It should be noted that this process is optional, as in many cases, the quality of the learned geometry and appearance without refinement is already satisfactory.
5 Experiments
In this section, we detail the implementation across different stages and describe the datasets used for evaluation.
@@ 5 Experiments
5.1 Implementation details
The proposed method is implemented using PyTorch and the rasterization toolkit from 3DGS [14]. Experiments are conducted on an RTX A6000 GPU with 24GB memory. We will release source code upon acceptance of the paper.
5.1.1 Initialization.
We train the original 3DGS model for 3,000 iterations to obtain a set of Gaussians, whose center positions are treated as a point cloud. We employ Alpha Shapes [4] to extract a coarse mesh and make it watertight [11]. In the next stage, we will construct an SDF grid based on this coarse mesh.
Our approach differs from others [19, 23] that randomly initialize the SDF, which would cause an OOM issue when the scene becomes large. Thanks to the fast optimization and the discrete nature of 3DGS, we are able to capture the coarse geometry within 5 minutes.
5.1.2 Joint Learning Mesh and Gaussians.
The resolution of the used SDF grid depends on the scene. For Mip-NeRF360 scenes, we use $150^{3}$ to $250^{3}$ grid cells for the foreground and $250^{3}$ to $300^{3}$ for the background, placing 3 Gaussians for every marched face. For NeRF-Synthetic scenes, we use $100^{3}$ grid cells and place 6 Gaussians on every face. For the appearance model, we adopt an MLP with 2 hidden layers, each comprising 32 neurons. This stage requires 10k iterations for optimization, totally using 15-50 minutes.
5.1.3 Refinement.
The adaptive Gaussian binding strategy and appearance model enable mesh modifications within a specified range without impacting performance. Thus, we perform subdivision or decimation on the mesh to unify the number of faces (0.1M for single objects, 1.8M for real-world scenes) and then re-calculate the Gaussian covariances and colors. The refinement process consists of 7k iterations and requires 15-50 minutes.
5.2 Datasets and Metrics
@@
Table 1: Per-scene quantitative comparisons on NeRF-Synthetic dataset [17].
@@
Ours 34.48 0.980 0.018 28.81 0.953 0.039 35.44 0.992 0.007 28.58 0.882 0.124
Table 2: Quantitative comparisons on Mip-NeRF360 dataset [1].
@@
Ours 29.06 0.908 0.198 23.80 0.665 0.274 26.80 0.804 0.230
Figure 4: Qualitative comparisons with SuGaR [10] on the Mip-NeRF360 dataset [1].
5.3 Novel View Synthesis
We first evaluate our method on the novel view synthesis task. Tabs. 1 and 2 present the quantitative results on NeRF-Synthetic and Mip-NeRF360 datasets. To ensure a fair comparison, we employed a white background for the NeRF-Synthetic dataset. Our approach is compared against various recent methods, including those with reconstructed mesh and those without.
For the NeRF-Synthetic dataset, our method achieves state-of-the-art rendering quality among methods reconstructing a mesh, and even outperforms some models that solely focus on rendering quality. Furthermore, there is only a slight decrease in performance compared to the original 3DGS. For the Mip-NeRF360 dataset, our method is comparable to the state-of-the-art method [10]. In some cases (e.g., bicycle), our method achieves better geometric correctness, even if the PSNR for those scenes is not as high as SuGaR’s. This phenomenon is demonstrated by the qualitative comparison in Fig. 4.
Figure 5: Rendered images and reconstructed meshes on NeRF-Synthetic [17] and Mip-NeRF360 [1] datasets.
Table 3: Mesh quality evaluation. Chamfer Distance (CD) are compared to the ground truth meshes from NeRF-Synthetic dataset.
Method CD ($10^{-3}$) #Faces PSNR
NVdiffRec [19] 8.15 80k 29.05
NeRF2Mesh [26] 5.06 192k 31.04
Ours 7.27 100k 32.32
5.4 Surface Reconstruction
Although our focus is on establishing a hybrid representation that ensures alignment between Gaussians and mesh faces, the surfaces reconstructed by our method are of good quality. We visualize the reconstructed meshes and their normal maps in Fig. 5, which demonstrates that our method effectively captures the surface details of a scene. More results can be found in the appendix.
For a quantitative evaluation of the meshes, we compare the reconstructed mesh to the ground truth from the NeRF-Synthetic dataset using the Chamfer Distance (CD) metric. The computation method follows NeRF2Mesh [26], where 2.5M points are sampled from the surfaces of both the ground truth and the reconstructed object through ray casting. Tab. 3 presents the CD results, the number of faces, and the rendering PSNR, illustrating that our resultant meshes achieve high quality in quantitative measures. Furthermore, the Gaussian-based appearance enables high-quality rendering even with fewer faces than NeRF2Mesh.
@@ 5.4 Surface Reconstruction
5.5 Object Deformation by Hybrid Representation
As discussed in our introduction section, the hybrid representation not only benefits from the high-quality and efficient rendering of Gaussians but also offers the convenience of manipulation through the mesh. To demonstrate the potential applications of this work, we selected object deformation tasks for testing. Specifically, we deform the learned mesh using four common mesh modifiers (Twist, Stretch, Bend, and Taper) in Blender and correspondingly adjust the bound Gaussians using the technique described in Sec. 4.2. We then use the Gaussian-based rasterization to render images of the deformed object, as shown in Fig. 6. More advanced operations, as explored in [12, 9], have been examined and are also compatible with the representation we have obtained.
5.6 Method Analysis
@@
Table 4: Ablation studies. We conduct evaluations on the Mip-NeRF360 dataset [1] and report the average PSNR across all scenes.
Adaptive Covariance Refinement PSNR
x $\checkmark$ 26.53
$\checkmark$ x 23.77
$\checkmark$ $\checkmark$ 26.80
GS/face PSNR Time
1 26.53 1.5h
3 26.80 2h
6 26.87 2.5h
5.6.1 Adaptability in Scene Updates.
One unique feature of our method is the simultaneous learning of topology (defined by mesh faces) and Gaussians. In contrast, the existing method, SuGaR [10], separates the learning of topology and Gaussians, as the mesh extraction process (Poisson Surface Reconstruction) is non-differentiable for Gaussian rasterization.
In practice, when part of the scene requires updating, our model can directly adapt to these changes without re-learning from scratch. Fig. 7 presents the results of dealing with scene updates using our method and SuGaR. Our method successfully updates the geometry and appearances of the updated regions, while SuGaR struggles to adapt to new scene geometries with its fixed topology.
5.6.2 Ablation Studies.
In this section, we examine the impact of three factors: 1) the adaptive Gaussian covariance; 2) the refinement stage; and 3) the number of Gaussians per face. For evaluating the adaptive Gaussian covariance, we set $\boldsymbol{M}$ in Eq. (9) to be an identity matrix.
Tab. 4 presents results from the Mip-NeRF360 dataset, showing that both adaptive Gaussian covariance and the refinement stage significantly enhance performance. Additionally, increasing the number of Gaussians per face is beneficial for rendering quality but leads to a longer training time. As a trade-off, we use only 3 Gaussians per face in this paper.
6 Conclusion

### v2

@@ IV-B Gaussians Consistent with Scene Geometry
Before determining the Gaussian covariances, we introduce a local coordinate frame as illustrated in Fig. 3 (b). This frame has its origin at $\boldsymbol{v}_{1}$. The axes of the frame are denoted as $\boldsymbol{t}_{i}=\tilde{\boldsymbol{t}}_{i}/\|\tilde{\boldsymbol{t}}_{i}\|$, $i\in\{1,2,3\}$, where
$\tilde{\boldsymbol{t}}_{1}=\boldsymbol{a}\times\boldsymbol{b},\quad\tilde{\boldsymbol{t}}_{2}=\boldsymbol{a},\quad\tilde{\boldsymbol{t}}_{3}=(\boldsymbol{a}\times\boldsymbol{b})\times\boldsymbol{a},$ (9)
@@
If the angles of a triangle are all close to $\pi/3$, the Gaussian covariances in the local coordinate frame can be naturally defined as $\boldsymbol{\Sigma}_{e}=\text{diag}(\epsilon,r^{2},r^{2})$, representing a thin layer of Gaussian surrounding a small area of the surface. The choice of $r$ is empirical. For triangles with a more irregular aspect ratio, where the area is challenging to approximate by a disc, it is beneficial to apply a linear transformation $\boldsymbol{M}$ on $\boldsymbol{\Sigma}_{e}$ as visualized in Fig. 3 (c). Hence, the covariance matrix in world coordinate system is given by:
@@
The derivation of matrix $\boldsymbol{M}$ can be found in the appendix.
IV-C Gaussian Appearance Model
In the original 3DGS model, the per-Gaussian colors $\{\boldsymbol{c}_{k}\}$ are optimized directly because the Gaussians remain consistent throughout the training process. However, in our framework, the Gaussians are dynamically constructed from the evolving SDF. Unlike scene geometry, where gradients from the Gaussian centers $\{\boldsymbol{\mu}_{k}\}$ flow back through the same parametric representation $\{s_{n}\}_{n=1}^{N}$ of the SDF, this dynamic construction makes continuous color optimization difficult.
Figure 4: Comparison between two approaches to learn appearance.
Specifically, the key challenges we face is the lack of direct connections between Gaussians derived from $(\mathcal{V}_{t},\mathcal{F}_{t})$ and those from $(\mathcal{V}_{t-1},\mathcal{F}_{t-1})$, where $t$ denotes the $t$-th training step. A straightforward approach to overcome this would be to compute colors based on their neighboring Gaussians from the previous step, formulated as $\boldsymbol{c}_{i,t}=\frac{1}{|\mathcal{N}|}\sum_{j\in\mathcal{N}}\boldsymbol{c}_{j,t-1}$, where $\boldsymbol{c}_{i,t}$ denotes the color of the $i$-th Gaussian at step $t$, and $\mathcal{N}$ is the set of neighbors based on the Gaussian centers. However, this approach presents two significant drawbacks: it requires substantial computation due to the neighbor search between two large sets of Gaussians, averaging colors within a local region might lead to low effectiveness in learning scene details, as shown in Fig. 4 (b).
To address these challenges, we introduce a neural appearance model, $\mathcal{A}:\mathbb{R}^{3}\mapsto\mathbb{R}^{d}$, with $d$ representing the number of the used Spherical Harmonics coefficients. At each step, we determine the color as $\boldsymbol{c}=\mathcal{A}(\boldsymbol{\mu})$ for the Gaussian centered at $\boldsymbol{\mu}$. Thus, the appearance of Gaussians depends on the learnable parameters within $\mathcal{A}$, which remain consistent during training.
We implement $\mathcal{A}$ using hash-grid positional encoding [18] and a lightweight MLP to enhance efficiency and achieve a compact and continuous representation, inspired by NvdiffRec [19]. This neural appearance model can be robustly learned, even as the mesh faces changes from $(\mathcal{V}_{t-1},\mathcal{F}_{t-1})$ to $(\mathcal{V}_{t},\mathcal{F}_{t})$. Fig. 4 illustrates the experimental comparison between the two approaches. The neural appearance model can achieve nearly 40x speed-up and better learned details than the straightforward nearest neighbor searching approach.
@@ IV-C Gaussian Appearance Model
IV-D Rendering and Optimization
As the Gaussians we construct retain the same form as those in the original 3DGS, we employ the same method defined in Eq. (3) for rendering an image. For scenes with a background, the rendered Gaussians include both the Gaussians derived from the mesh and the additional background Gaussians. Our method relies only on photometric supervision and utilizes the loss function:
@@ IV-D Rendering and Optimization
Upon the convergence of the mesh and appearance model, we fix the topology defined by faces $\mathcal{F}$ and proceed to jointly refine the vertices and the associated 3D Gaussians. During this phase, Gaussian centers continue to be derived from the face vertices using established barycentric coordinates. However, for covariance and color, we adopt a different approach since we can optimize a consistent set of Gaussian attributes during this phase. Specifically, we sample the spherical harmonics (SHs) from the appearance model $\mathcal{A}$ and make them a learnable, per-Gaussian variable as in the original 3DGS. Moreover, the Gaussian covariance is parameterized into learnable 2D scaling $\hat{\boldsymbol{s}}\in\mathbb{R}^{2}$ and a complex number $\boldsymbol{a}i+\boldsymbol{b}$ representing rotation, which are similar to the process in SuGaR [10]. It should be noted that this process is optional, as in many cases, the quality of the learned geometry and appearance without refinement is already satisfactory.
V Experiments
In this section, we first detail the implementation across different stages and describe the datasets used for evaluation.
@@ V Experiments
V-A Implementation details
The proposed method is implemented using PyTorch and the rasterization toolkit from 3DGS [14]. Experiments are conducted on an RTX3090 GPU with 24GB memory.
V-A1 Initialization
We train the original 3DGS model for 20,000 iterations to obtain a set of Gaussians, whose center positions are treated as a point cloud. We then employ Alpha Shapes [4] to extract a coarse mesh and make it watertight [11], which is subsequently used to construct an SDF grid. Additionally, the background Gaussians are retained and continuously trained in the later stages.
Our approach differs from others [19, 23] that randomly initialize the SDF, which would cause an OOM issue when the scene becomes large. Thanks to the fast optimization and the discrete nature of 3DGS, we are able to capture the coarse geometry using only 5-20 minutes.
V-A2 Joint Learning of Mesh and Gaussians
The resolution of the used SDF grid depends on the scene. For Mip-NeRF360 scenes, we use $150^{3}$ to $250^{3}$ grid cells for the foreground, attaching 3 Gaussians to every face. For NeRF-Synthetic scenes, we use $100^{3}$ grid cells and attach 3 Gaussians to every face. For the appearance model, we adopt an MLP with 2 hidden layers, each comprising 32 neurons. This stage requires 10k iterations for optimization, totally using 15-30 minutes.
V-A3 Refinement
Before the refinement stage, we perform subdivision and decimation on the mesh to unify the number of faces (0.1M for single objects, 0.5M for real-world scenes). We attach 6 Gaussians to each face and compute their covariances and colors using the adaptive Gaussian construction and appearance model, with minimal impact on model performance. The refinement process consists of 5-10k iterations and requires 10-20 minutes.
V-B Datasets and Metrics
@@
TABLE I: Per-scene quantitative comparisons on NeRF-Synthetic dataset [17].
@@
SuGaR-15K [10] 34.87 0.981 0.015 27.86 0.944 0.046 34.81 0.989 0.008 29.22 0.880 0.103
Ours 34.94 0.982 0.016 27.93 0.950 0.043 34.89 0.991 0.007 28.32 0.883 0.121
TABLE II: Quantitative comparisons on Mip-NeRF360 dataset [1].
@@
Ours 29.33 0.910 0.193 25.17 0.754 0.237 27.54 0.843 0.212
Figure 5: Qualitative comparisons with baseline methods (SuGaR [10], NeRF2Mesh [26]) on the NeRF-Synthetic [17] and Mip-NeRF360 dataset [1].
V-C Novel View Synthesis
We first evaluate our method on the novel view synthesis task. Tabs. I and II present the quantitative results on NeRF-Synthetic and Mip-NeRF360 datasets. To ensure a fair comparison, we employed a white background for the NeRF-Synthetic dataset. Our approach is compared against various recent methods, including those with reconstructed mesh and those without.
For the NeRF-Synthetic dataset, our method achieves state-of-the-art rendering quality among methods reconstructing a mesh, and even outperforms some models that solely focus on rendering quality. Furthermore, there is only a slight decrease in performance compared to the original 3DGS [14]. For the Mip-NeRF360 dataset, our method is comparable to or surpasses the state-of-the-art methods [10, 26]. Fig. 5 includes qualitative comparisons between images rendered by our method and the baselines, demonstrating that our method captures better scene details.
Figure 6: Qualitative comparisons of reconstructed meshes with SuGaR [10] on the NeRF-Synthetic [17] and Mip-NeRF360 [1] datasets.
TABLE III: Quantitative Evaluation of Mesh Quality and Efficiency on the NeRF-Synthetic [17] dataset.
Method CD ($10^{-3}$) #Faces PSNR Train Time
NVdiffRec [19] 8.15 80k 29.05 52 mins
NeRF2Mesh [26] 5.06 192k 31.04 46 mins
SuGaR-15K [10] 8.62 1000k 32.09 103 mins
Ours (w/o refine) 8.50 141k 27.68 16 mins
Ours 7.27 100k 32.26 28 mins
‘CD’: Chamfer Distance between the ground truth and reconstructed meshes.
V-D Surface Reconstruction
Although our focus is on efficiently learning a hybrid representation that ensures alignment between Gaussians and mesh faces, the surfaces reconstructed by our method are of high quality. In Fig. 6, we visualize the reconstructed meshes and their normal maps, comparing our method against the baseline. The results demonstrate that our method effectively captures the surface details of a scene, whereas SuGaR [10] may produce floating patches and holes in the mesh.
For a quantitative evaluation of the meshes, we compare the reconstructed mesh to the ground truth from the NeRF-Synthetic dataset using the Chamfer Distance (CD) metric. The computation method follows NeRF2Mesh [26], where 2.5M points are sampled from the surfaces of both the ground truth and the reconstructed object through ray casting. For a more comprehensive assessment, Tab. III presents the CD results, the number of faces, the rendering PSNR, and the training time. The results demonstrate that our reconstructed meshes achieve both high quality and training efficiency. Additionally, our Gaussian-based appearance model enables high-quality rendering even with fewer faces compared to NeRF2Mesh.
Furthermore, the learned mesh without refinement is already of satisfactory quality. This indicates that for applications requiring only an untextured mesh, our method can produce excellent results very efficiently, with a training time of just 16 minutes.
@@ V-D Surface Reconstruction
V-E Object Deformation by Hybrid Representation
As discussed in our introduction section, the hybrid representation not only benefits from the high-quality and efficient rendering of Gaussians but also offers the convenience of manipulation through the mesh. To demonstrate the potential applications of this work, we selected object deformation tasks for testing. Specifically, we deform the learned mesh using four common mesh modifiers (Twist, Stretch, Bend, and Taper) in Blender and correspondingly adjust the bound Gaussians using the technique described in Sec. IV-B. We then use the Gaussian-based rasterization to render images of the deformed object, as shown in Fig. 7. More advanced operations, as explored in [12, 9], are also compatible with the representation we have obtained.
V-F Method Analysis
@@
TABLE IV: Evaluation on Efficiency. We report the training time and PSNR metric across different methods.
Method NeRF-Synthetic [17] Mip-NeRF360 [1]
[…]
NVdiffRec [19] 29.05 52 mins - -
NeRF2Mesh [26] 31.04 46 mins - -
SuGaR-15K [10] 32.09 103 mins 27.27 132 mins
Ours 32.26 28 mins 27.54 61 mins
‘-’ indicates that the method does not have results reported in their paper and fails to reconstruct all scenes of the dataset using their published codes.
TABLE V: Ablation studies on adaptive covariance and refinement stage. We report the average PSNR across all scenes of the NeRF-Synthetic [17] and Mip-NeRF360 [1] datasets.
Adaptive Cov. Refine NeRF-Synthetic [17] Mip-NeRF360 [1]
x $\checkmark$ 32.02 27.19
$\checkmark$ x 27.78 24.44
$\checkmark$ $\checkmark$ 32.26 27.54
TABLE VI: Ablation Study on the Number of Gaussians per Face. This table shows the PSNR results on the Mip-NeRF360 [1] dataset.
PSNR Stage-2
#.GS/face=1 #.GS/face=3 #.GS/face=6
Stage-1 #.GS/face=1 26.80 27.12 27.22
#.GS/face=3 26.89 27.21 27.54
‘Stage-1’: Joint learning of mesh and Gaussian appearance.
‘Stage-2’: Refinement stage that fixes mesh faces.
Note: The table does not include #.GS/$\geq$6 for Stage-1 due to OOM issues in some scenes.
V-F1 Efficiency
We assess our method’s efficiency by comparing the average time taken to reconstruct scenes from the NeRF-Synthetic and Mip-NeRF360 datasets. Both our method and the baseline methods are tested under identical hardware conditions. The results in Tab. IV reveal that our method is not only faster but also provides superior rendering quality.
V-F2 Adaptability in Scene Updates
One unique feature of our method is the simultaneous learning of topology (defined by mesh faces) and Gaussians. In contrast, the existing method, SuGaR [10, 27, 9], separates the learning of topology and Gaussians, as the mesh extraction process (Poisson Surface Reconstruction) is non-differentiable for Gaussian rasterization. In practice, when part of the scene requires updating, our model can directly adapt to these changes without re-learning from scratch, whereas the existing method cannot. Fig. 8 presents the results of handling scene updates using our method and SuGaR. Our method successfully updates the geometry and appearances of the modified regions, while SuGaR struggles to adapt to new scene geometries with its fixed topology.
V-F3 Ablation Studies
In this section, we examine the impact of three factors: 1) the adaptive Gaussian covariance, 2) the refinement stage, and 3) the number of Gaussians per face. To evaluate the effect of adaptive Gaussian covariance, we set $\boldsymbol{M}$ in Eq. (10) to be an identity matrix. Tab. V shows that both adaptive Gaussian covariance and the refinement stage significantly enhance rendering quality.
The choice of the number of Gaussians per face involves a trade-off between computational consumption and performance. As shown in Tab. VI, increasing the number of Gaussians per face benefits rendering quality. To control the memory footprint and training time within an acceptable range while still achieving excellent rendering quality, we choose 3 Gaussians per face during the joint learning of mesh and Gaussian appearance, but use 6 Gaussians for the refinement stage.
VI Conclusion